OpenStudy (anonymous):

Factor completely and check mentally 8x^2yz+16xy^2z^2-24xyz^2

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

a term is separated by a + or - sign

OpenStudy (anonymous):

xyz

OpenStudy (dan815):

tell me what numbers or variables are common in all those terms, there are 3 terms in here

OpenStudy (dan815):

good, what about any numbers

OpenStudy (anonymous):

No numbers.

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

8 16 and 24

OpenStudy (dan815):

remember a number can be a multiple

OpenStudy (dan815):

right remember 16 = 8*2 and 24=8*3

OpenStudy (dan815):

okay so lets factor 8xyz(_____ + _______ - _______) <--- what goes in the blanks

OpenStudy (dan815):

think what do i have to multiply to get back the terms you started with

OpenStudy (anonymous):

16xyz and 24xyz?

OpenStudy (dan815):

ill do the first term for you okay

OpenStudy (anonymous):

._. Alright.

OpenStudy (dan815):

we are factoring 8xyz and we know our first term is 8x^2yz, so that means i only need to multiply 8xyz*x to get back what i started with so... first blank is x, do the other blanks 8xyz(___x___+ _______ - _______) <--- what goes in the blanks

OpenStudy (anonymous):

Oh hey if I get a bit rushy please disregard class has been over for 7 minutes and I'm on question 4/10

OpenStudy (dan815):

I told you I am worse than mathmale, figure this out now

OpenStudy (dan815):

This is as much as I can help you

OpenStudy (anonymous):

Hold on.. x? Why x?

OpenStudy (muzzack):

Step by step solution : Step 1: Simplify \(8x^2yz+16xy^2z^2 - (2^3•3xyz^2)\) Pull out like factors : \(8x^2yz + 16xy^2z^2 - 24xyz^2 = 8xyz • (x + 2yz - 3z)\) Final result : \(\huge\ 8xyz • (x + 2yz - 3z)\)